OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − 2y = 2 3x + y = 6

OpenStudy (anonymous):

The first equation can be turned into x=2+2y so then substitute 2+2y into the second equation and solve.

OpenStudy (anonymous):

so 3(2+2y)=6-y?

OpenStudy (wach):

An easy way to find x is to completely eliminate y using systems of equations. Is there any way to manipulate the equations so that when added, y is eliminated?

OpenStudy (amistre64):

flow, your substitution is fine

OpenStudy (campbell_st):

to eliminate y double the 2nd equation and add to the 1st eauation x - 2y = 2 6x + 2y = 12 ---------------------- 7x = 14 solve for x then substitute to find y

OpenStudy (amistre64):

but to find "x", you want to define y(x) for a sunstitution

OpenStudy (amistre64):

x − 2y = 2 3x + y = 6 <-- simpler to use y(x) = 6-3x x − 2y(x) = 2 x − 2(6-3x) = 2